OpenStudy (anonymous):

C=0.56(F-32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 8 degrees Celsius is.

OpenStudy (anonymous):

that is the value in C of 8F, part 2 asks for the reverse :-) for part 1, expand the right hand side then isolate F. Let me know if you want a hand getting started.

OpenStudy (anonymous):

Part 2: 8º Celsius = 0.56(F-32) = 46.4º Fahrenheit

OpenStudy (radar):

C=0.56(F-32) Divide both sides by 0.56 getting: C/0.56 = F-32 Add 32 to both sides, getting: (C/0.56) + 32 = F or F=(C/0.56) + 32 That is part 1. Use that formula to do convert the 8 degree Celsius.

OpenStudy (radar):

Part 2: F=(8/0.56)+32 F=46.3 degrees

OpenStudy (radar):

Sorry but my computer got hung up when "owl be back" came on lol